40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 25 23:20:57 2013, Annalisa, Summary, Auxiliary locking, Beat note found! 
    Reply  Thu Mar 28 19:32:22 2013, Annalisa, Summary, Auxiliary locking, Beat note found! BeatFreq.jpg
Message ID: 8345     Entry time: Mon Mar 25 23:20:57 2013     Reply to this: 8368
Author: Annalisa 
Type: Summary 
Category: Auxiliary locking 
Subject: Beat note found! 

[Annalisa, Manasa]

The beat note between the main PSL and the auxiliarly NPRO has been found!

The setup didn't change with respect to the one described on the previous note on the elog. A multimeter has been connected to the laser controller diagnostic pin to read out the voltage that indicated the laser crystal temperature.

The connector has been taken from the Yend table laser controller.

The voltage on the multimeter gave the same temperature shown by "Laser temperature" on the display of the controller, while "set temperature" was wrong.

The temperature has been varied using the laser controller with reference to the voltage read on the multimeter display.

Starting from 35.2 °C, the temperature has been first lowered until 20 °C and no beat note has been found, then temperature has been increased up to 35.2 °C and the first beat note has been found at 38.0 °C.

It has been detected at a frequency of about 80 MHz with an RF power of -27 dBm and a frequency fluctuation of about  +/- 4 MHz.

To do:

I made more measurements slowly varying the laser temperature, to see how the beat note frequency changes with it. I'll make the plot and post it as soon.

ELOG V3.1.3-